### v1

@@ Abstract
Context. Dark matter (DM) self-interactions alter the matter distribution on galactic scales and alleviate tensions with observations. A feature of the self-interaction cross section is its angular dependence, influencing offsets between galaxies and DM halos in merging galaxy clusters. While algorithms for modelling mostly forward-dominated or mostly large-angle scatterings exist, incorporating realistic angular dependencies within $N$-body simulations remains challenging.
Aims. To efficiently simulate models with a realistic angle-dependence, such as light mediator models, we develop, validate and apply a novel method.
Methods. Existing approaches to describe small- and large-angle scattering regimes are combined within a hybrid scheme. Below a critical angle the effective description of small-angle scattering via a drag force combined with transverse momentum diffusion is used, while above the angle-dependence is sampled explicitly.
Results. First, we verify the scheme using a test set-up with known analytical solutions, and check that our results are insensitive to the choice of the critical angle within an expected range. Next, we demonstrate that our scheme speeds up the computations by multiple orders of magnitude for realistic light mediator models. Finally, we apply the method to galaxy cluster mergers and discuss the sensitivity of the offset between galaxies and DM to the angle-dependence of the cross section. Our scheme ensures accurate offsets for mediator mass $m_{\phi}$ and DM mass $m_{\chi}$ within the range $0.1v/c\lesssim m_{\phi}/m_{\chi}\lesssim v/c$, while for larger (smaller) mass ratios the offsets obtained for isotropic (forward-dominated) self-scattering are approached. Here $v$ is the typical velocity scale. Equivalently, the upper condition can be expressed as $1.1\lesssim\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\lesssim 10$ for the ratio of total and momentum transfer cross sections, with the ratio being $1$ ($\infty$) in the isotropic (forward-dominated) limits.
@@ 1 Introduction
While astrophysical observations ranging from galactic to cosmological scales support the existence of dark matter (DM) via its gravitational impact, it is plausible that DM features also non-gravitational interactions, like any other measured particle species. Extensive searches for non-gravitational interactions between dark and visible matter have not found an unambiguous signal so far, and pushed the limits even beyond the scale of weak interactions in many cases. In contrast, interactions of DM with itself with sizeable strength, comparable to that known from the strong interaction, remain a viable possibility, that is testable by searching for its imprints on the dynamics of galaxies and galaxy clusters.
@@
Elastic $2\to 2$ DM self-interactions have been proposed by Spergel & Steinhardt (2000) in the context of the core-cusp problem, with early $N$-body simulations by Burkert (2000) showing core formation followed by gravothermal collapse. Self-interacting dark matter (SIDM) has also been proposed for addressing other small-scale puzzles, such as the abundance and properties of satellite galaxy populations or the variability observed in galaxy rotation curves, see e.g. Bullock & Boylan-Kolchin (2017); Tulin & Yu (2018); Adhikari et al. (2022) for reviews.
@@
While SIDM on galactic scales typically requires a cross section (per DM mass) of the order of $\sigma/m\gtrsim 1\text{cm}^{2}\text{g}^{-1}$ to have a sizeable impact, galaxy cluster observations have been argued to require upper limits reaching down to $0.1\text{cm}^{2}\text{g}^{-1}$ (see e.g. Elbert et al., 2018; Sagunski et al., 2021; Andrade et al., 2021; Despali et al., 2022; Eckert et al., 2022; Zhang et al., 2024). Given the very different typical velocities on cluster and galactic scales, this suggests a velocity-dependent cross section (see for example Kaplinghat et al., 2016), such as occurs e.g. if the interaction is mediated by an exchange particle with mass $m_{\phi}$ that is light compared to the DM mass $m_{\chi}$ (Feng et al., 2009; Buckley & Fox, 2010; Loeb & Weiner, 2011; Tulin et al., 2013).
Within this class of models, the scattering cross section also features a pronounced angular dependence, being strongly forward-dominated in the limit $m_{\phi}\to 0$, analogously to Rutherford scattering.
@@
For isolated, quasi-stationary halos the impact of angular dependence on the density profile has been argued to be approximately captured by mapping simulations assuming isotropic scattering onto models with anisotropic differential cross sections by matching a suitable angle-averaged effective cross section, namely the viscosity cross section (e.g. Colquhoun et al., 2021; Yang & Yu, 2022; Sabarish et al., 2024).
However, for systems not being quasi-stationary, such as for example the evolution of satellite populations (Kahlhoefer et al., 2015; Fischer et al., 2022, 2024b; Ragagnin, A. et al., 2024), when approaching gravothermal collapse, and especially for strongly non-stationary processes such as for example galaxy cluster mergers (e.g. Kahlhoefer et al., 2014; Robertson et al., 2017a; Fischer et al., 2021a, b, 2023; Sabarish et al., 2024), it is a priori less clear how they are affected by (strongly) anisotropic scattering.
@@
This motivates to develop efficient numerical algorithms for incorporating differential cross sections featuring a pronounced angular dependence, such as for light mediator models, in order to assess the accuracy of (and possibly refine) commonly employed approximate ‘mapping’ prescriptions for a given observable, and potentially identify new features that are specific to a given angular dependence. In this work we develop such an algorithm, validate it and apply it to study galaxy cluster mergers with angle-dependent cross sections that are characteristic for light mediator models.
@@
Our work is based on a combination of two existing approaches, one that is suitable for relatively rare individual self-scattering events, each of which features large deflection angles (dubbed rare SIDM or ‘rSIDM’, see Sect. 2 and e.g. Koda & Shapiro (2011); Vogelsberger et al. (2012); Rocha et al. (2013); Fry et al. (2015); Robertson et al. (2017a); Banerjee et al. (2020) for details). In this case SIDM can be modelled by collisions of $N$-body particles, with scattering angles sampled from the differential cross section. While being most straightforward and direct, it has the drawback of becoming numerically prohibitively inefficient for models in which the scattering rate is large but the momentum transfer in each individual scattering event is small.
@@
The second approach is tailored precisely to efficiently describe such very frequent and strongly forward dominated scatterings (dubbed frequent SIDM or ‘fSIDM’). The fSIDM approach formally corresponds to the limit in which the momentum transfer cross section is kept constant but the typical scattering angle approaches zero. Such frequent interactions are well-known in the context of Coulomb interactions of energetic particles traversing a medium, and can effectively be captured by a drag force complemented with transverse momentum diffusion. A drag force description had been derived by Kahlhoefer et al. (2014) and applied to merging galaxy clusters. Building up on that work, an algorithm for the fSIDM limit has been proposed in the context of $N$-body simulations of SIDM by Fischer et al. (2021a) (respecting energy and momentum conservation). Their implementation has been developed further by Fischer et al. (2021b, 2022, 2024b) and used by Fischer et al. (2023, 2024a); Fischer & Sagunski (2024); Sabarish et al. (2024); Ragagnin, A. et al. (2024) to study the SIDM halo evolution, galaxy cluster merger dynamics, the evolution of cluster satellites and dynamical friction as compared to the usually considered opposite limit of isotropic rSIDM. However, as mentioned above, realistic models feature both small- and large-angle scattering with roughly comparable impact over a wide range of parameter space, and thus neither the fSIDM nor rSIDM scheme can be considered to describe viable particle models in general.
@@
In this work we therefore develop a new hybrid method, that we dub ‘hSIDM’ scheme, that is designed to be able to accurately describe (in principle) arbitrary differential cross sections, featuring in particular a strongly forward dominated enhancement at low deflections angles as well as a given contribution with large-angle scattering. The hSIDM method is therefore particularly suited to study light mediator models, but can easily be extended to other scenarios that have been proposed in the literature, see e.g. Tulin & Yu (2018) for an overview. The hSIDM algorithm takes advantage of the numerically much more efficient fSIDM approach for small scattering angles, below a certain critical angle $\theta_{\mathrm{c}}$, complemented by the more conventional rSIDM treatment of large-angle scatterings.
@@
Having developed the hSIDM scheme we apply it to merging galaxy clusters. Systems such as the ‘Bullet Cluster’ (e.g. Springel & Farrar, 2007; Mastropietro & Burkert, 2008; Lage & Farrar, 2014) or ‘El Gordo’ (e.g. Donnert, 2014; Molnar & Broadhurst, 2015; Zhang et al., 2015; Kim et al., 2021; Asencio et al., 2020, 2023) have been studied and used to constrain DM self-interactions (e.g. Valdarnini, R., 2024). Besides other phenomena, such as the oscillations of the brightest cluster galaxy (BCG) at the late stages of the merger (e.g. Harvey et al., 2019; Cross et al., 2024), offsets between the distribution of the DM and the cluster galaxies mainly shortly after the first pericentre passage received a lot of attention (e.g. Harvey et al., 2015; Robertson et al., 2017b; Sirks et al., 2024). Theoretical studies to model those offsets that arise from the self-interactions effectively decelerating the two DM halos when they pass through each other while the galaxies are only indirectly affected via gravity have been conducted to gain insights into their evolution and allow for inferring constraints from observed systems. Offsets between the galaxy and DM distributions are measured employing strong gravitational lensing to infer the total mass distribution. These measurements are challenging and previous claims of large offsets have been questioned (e.g. Bradač et al., 2008; Dawson et al., 2012; Dawson, 2013; Jee et al., 2014, 2015; Harvey et al., 2017; Peel et al., 2017; Taylor et al., 2017; Wittman et al., 2018, 2023). Until today observed systems show no clear deviation from collisionless DM and mainly upper bounds on the self-interaction cross section from DM-galaxy offsets exist.
@@
This work is structured as follows. In Sect. 2 we first introduce the various exemplary differential cross sections considered in this work and review the rSIDM and fSIDM schemes for describing angle-dependent self-scattering. We introduce the new hSIDM scheme in Sect. 3 and furthermore discuss an analytically solvable test set-up that we use for validation later on. The numerical implementation is discussed in Sect. 4, and we present our validation checks in Sect. 5. The hSIDM scheme is used to simulate galaxy cluster mergers for SIDM with angle-dependence as predicted by light mediator models in Sect. 6. Our main results for DM-galaxy offsets and their model-dependence is discussed are Sect. 7. Finally, we summarise and conclude in Sect. 8. Additional information can be found in the appendices.
2 Review of angle-dependence in SIDM
@@
In this section, we first briefly review the angle-dependence of typical differential cross sections relevant for light mediator models of SIDM as well as various definitions of angle-averaged cross sections considered in this context. We then review the existing rSIDM and fSIDM approaches for describing rare large-angle and frequent small-angle scatterings, respectively, on which the hybrid scheme introduced in this work is based.
@@ 2.1 Differential scattering cross sections
Light mediator models are characterised by scattering via a Yukawa interaction, for which DM particles of mass $m_{\chi}$ scatter via the exchange of a mediator with mass $m_{\phi}$. One can further discriminate the cases in which the two incoming DM particles are identical (we refer to this case as ‘Møller scattering’ in analogy to $e^{-}e^{-}$ scattering, noting however that the mediator particle is massive here), or distinguishable (being e.g. a particle and an antiparticle, referred to as ‘Rutherford scattering’, but again representing the case of a massive mediator in this work). In the non-relativistic limit and in Born approximation, Møller scattering proceeds via either a $t$- or $u$-channel diagram, yielding
@@
with the factor $1/2$ accounting for the double counting of the identical particles, while only the $t$-channel process contributes to Rutherford scattering, giving
@@
Following the notations used in Girmohanta & Shrock (2022); Robertson et al. (2017a), we introduced a parameter $\sigma_{0}$ that is related to the Yukawa coupling strength $\alpha_{\chi}$ via $\sigma_{0}=\alpha_{\chi}^{2}m_{\chi}^{2}/m_{\phi}^{4}$. More importantly, the angle-dependence is related to the parameter
@@
where $\beta_{\mathrm{rel}}=v_{\mathrm{rel}}/c$ is the relative velocity of the incoming particles. The value of $r$ characterises the dependence of the cross section on the deflection angle $\theta$, with the isotropic limit being approached for $r\ll 1$, while scattering becomes strongly forward-dominated (Coulomb-like) for large values of $r$. For illustration, we show the angle-dependence of the Møller cross section in Fig. 1 for $r=100$ and $10^{4}$, respectively, as well as an isotropic cross section for comparison. We emphasise the pronounced peaks towards small scattering angles (note the logarithmic $y$-axis). Since the Møller cross section is symmetric under $\theta\to\uppi-\theta$, it is also enhanced for $\theta\to\uppi$, due to the fact that both scattering particles are indistinguishable.
@@
Figure 1: Illustration of the differential cross section for self-scattering of identical particles via a light mediator, referred to as Møller scattering in this work, for two values of the anisotropy parameter $r$ defined in Eq. (3) (solid lines). Noting the logarithmic $y$-axis, the strong enhancement for $\theta\to 0$ and $\theta\to\uppi$ can be seen, that becomes the more pronounced the larger $r$ (i.e. the lighter the mediator mass). We note that the isotropic case (red dashed) is not flat due to the scaling of the cross section with $\sin(\theta)$ (such that the the lines represent the scattering rate within a given range $(\theta,\theta+\mathrm{d}\theta)$ of the azimuthal angle, for any value of the polar angle). We also illustrate the division into small and large-angle scattering by a critical angle $\theta_{\mathrm{c}}$, for two representative values, exemplifying the hybrid approach studied in this work.

### v2

@@ Abstract
Context. Dark matter (DM) self-interactions alter matter distribution on galactic scales and alleviate tensions with observations. A feature of the self-interaction cross section is its angular dependence, which influences offsets between galaxies and DM halos in merging galaxy clusters. While algorithms for modelling mostly forward-dominated or mostly large-angle scatterings exist, incorporating realistic angular dependencies within $N$-body simulations remains challenging.
Aims. To efficiently simulate models with a realistic angle dependence, such as light mediator models, we developed, validated, and applied a novel method.
Methods. We combined existing approaches to describe small- and large-angle scattering regimes within a hybrid scheme. Below a critical angle, the scheme uses the effective description of small-angle scattering via a drag force combined with transverse momentum diffusion, while above the angle, it samples the dependence explicitly.
Results. We first verified the scheme using a test set-up with known analytical solutions, and we checked that our results are insensitive to the choice of the critical angle within an expected range. Next, we demonstrated that our scheme speeds up the computations by multiple orders of magnitude for realistic light mediator models. Finally, we applied the method to galaxy cluster mergers. We discuss the sensitivity of the offset between galaxies and DM to the angle dependence of the cross section. Our scheme ensures accurate offsets for mediator mass $m_{\phi}$ and DM mass $m_{\chi}$ within the range $0.1v/c\lesssim m_{\phi}/m_{\chi}\lesssim v/c$, while for larger (smaller) mass ratios, the offsets obtained for isotropic (forward-dominated) self-scattering are approached. Here, $v$ is the typical velocity scale. Equivalently, the upper condition can be expressed as $1.1\lesssim\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\lesssim 10$ for the ratio of the total and momentum transfer cross sections, with the ratio being $1$ ($\infty$) in the isotropic (forward-dominated) limits.
@@ 1 Introduction
While astrophysical observations ranging from galactic to cosmological scales support the existence of dark matter (DM) via its gravitational impact, it is plausible that DM also features non-gravitational interactions as any other measured particle species does. Extensive searches for non-gravitational interactions between DM and visible matter have not found an unambiguous signal so far, and such work has pushed the limits even beyond the scale of weak interactions in many cases. In contrast, interactions of DM with itself with a sizeable strength, comparable to that known from the strong interaction, remain a viable possibility that is testable by searching for its imprints on the dynamics of galaxies and galaxy clusters.
@@
Elastic $2\to 2$ DM self-interactions have been proposed by Spergel & Steinhardt (2000) in the context of the core-cusp problem, with early $N$-body simulations by Burkert (2000) showing core formation followed by gravothermal collapse. Self-interacting DM (SIDM) has also been proposed for addressing other small-scale puzzles, such as the abundance and properties of satellite galaxy populations or the variability observed in galaxy rotation curves (see Bullock & Boylan-Kolchin (2017); Tulin & Yu (2018); Adhikari et al. (2022) for reviews).
@@
While SIDM on galactic scales typically requires a cross section (per DM mass) of the order of $\sigma/m\gtrsim 1\text{cm}^{2}\text{g}^{-1}$ to have a sizeable impact, galaxy cluster observations have been argued to require upper limits reaching down to $0.1\text{cm}^{2}\text{g}^{-1}$ (see e.g. Elbert et al., 2018; Sagunski et al., 2021; Andrade et al., 2021; Despali et al., 2022; Eckert et al., 2022; Zhang et al., 2024). Given the very different typical velocities on cluster and galactic scales, this suggests a velocity-dependent cross section (see for example Kaplinghat et al., 2016) such as occurs, for example, if the interaction is mediated by an exchange particle with mass $m_{\phi}$ that is light compared to the DM mass $m_{\chi}$ (Feng et al., 2009; Buckley & Fox, 2010; Loeb & Weiner, 2011; Tulin et al., 2013). Within this class of models, the scattering cross section also features a pronounced angular dependence, being strongly forward-dominated in the limit $m_{\phi}\to 0$, analogous to Rutherford scattering.
@@
For isolated quasi-stationary halos, the impact of angular dependence on the density profile has been argued to be approximately captured by mapping simulations assuming isotropic scattering onto models with anisotropic differential cross sections by matching a suitable angle-averaged effective cross section, namely the viscosity cross section (e.g. Colquhoun et al., 2021; Yang & Yu, 2022; Sabarish et al., 2024). However, for systems that are not quasi-stationary, such as during the evolution of satellite populations (Kahlhoefer et al., 2015; Fischer et al., 2022, 2024b; Ragagnin, A. et al., 2024) or when approaching gravothermal collapse, and especially for strongly non-stationary processes, such as galaxy cluster mergers (e.g. Kahlhoefer et al., 2014; Robertson et al., 2017a; Fischer et al., 2021a, b, 2023; Sabarish et al., 2024), it is a priori less clear how they are affected by (strongly) anisotropic scattering.
@@
These reasons motivated us to develop efficient numerical algorithms for incorporating differential cross sections featuring a pronounced angular dependence, such as for light mediator models, in order to assess the accuracy of (and possibly refine) commonly employed approximate ‘mapping’ prescriptions for a given observable and potentially identify new features that are specific to a given angular dependence. In this work, we develop such an algorithm, validate it, and apply it to study galaxy cluster mergers with angle-dependent cross sections that are characteristic for light mediator models.
@@
Our work is based on a combination of two existing approaches, including one that is suitable for relatively rare individual self-scattering events, which feature large deflection angles (dubbed rare SIDM or ‘rSIDM’; see Sect. 2 and e.g. Koda & Shapiro, 2011; Vogelsberger et al., 2012; Rocha et al., 2013; Fry et al., 2015; Robertson et al., 2017a; Banerjee et al., 2020, for details). In this first approach, SIDM can be modelled by collisions of $N$-body particles, with scattering angles sampled from the differential cross section. While being the most straightforward and direct, this method has the drawback of becoming prohibitively inefficient numerically for models in which the scattering rate is large but the momentum transfer in each individual scattering event is small.
@@
The second approach is tailored precisely to efficiently describe such very frequent and strongly forward-dominated scatterings (dubbed frequent SIDM or ‘fSIDM’). The fSIDM approach formally corresponds to the limit in which the momentum transfer cross section is kept constant but the typical scattering angle approaches zero. Such frequent interactions are well known in the context of Coulomb interactions of energetic particles traversing a medium and can effectively be captured by a drag force complemented with transverse momentum diffusion. A drag force description has been derived by Kahlhoefer et al. (2014) and applied to merging galaxy clusters. Building on that work, an algorithm for the fSIDM limit has been proposed in the context of $N$-body simulations of SIDM by Fischer et al. (2021a) (respecting energy and momentum conservation). Their implementation has been developed further by Fischer et al. (2021b, 2022, 2024b) and used by Fischer et al. (2023, 2024a); Fischer & Sagunski (2024); Sabarish et al. (2024); Ragagnin, A. et al. (2024) to study the SIDM halo evolution, galaxy cluster merger dynamics, the evolution of cluster satellites, and dynamical friction as compared to the usually considered opposite limit of isotropic rSIDM. However, as mentioned above, realistic models feature both small- and large-angle scattering with a roughly comparable impact over a wide range of parameter space, and thus neither the fSIDM nor rSIDM scheme can be considered to describe viable particle models in general.
@@
Therefore, in this work we develop a new hybrid method that we dub the ‘hSIDM’ scheme, and it is designed to be able to accurately describe (in principle) arbitrary differential cross sections. In particular, the hSIDM method can be used to efficiently simulate models featuring a strongly forward-dominated enhancement at low deflections angles as well as a given contribution with large-angle scattering. The hSIDM method is therefore particularly suited to study light mediator models, but it can easily be extended to other scenarios that have been proposed in the literature (see e.g. Tulin & Yu, 2018, for an overview). The hSIDM algorithm takes advantage of the numerically much more efficient fSIDM approach for small scattering angles below a certain critical angle $\theta_{\mathrm{c}}$ and is complemented by the more conventional rSIDM treatment of large-angle scatterings.
@@
After having developed the hSIDM scheme, we applied it to merging galaxy clusters. Systems such as the ‘Bullet Cluster’ (e.g. Springel & Farrar, 2007; Mastropietro & Burkert, 2008; Lage & Farrar, 2014) or ‘El Gordo’ (e.g. Donnert, 2014; Molnar & Broadhurst, 2015; Zhang et al., 2015; Kim et al., 2021; Asencio et al., 2020, 2023) have been studied and used to constrain DM self-interactions (e.g. Valdarnini, R., 2024). Besides other phenomena, such as the oscillations of the brightest cluster galaxy (BCG) at the late stages of the merger (e.g. Harvey et al., 2019; Cross et al., 2024), offsets between the distribution of the DM and the cluster galaxies mainly shortly after the first pericentre passage have received a lot of attention (e.g. Harvey et al., 2015; Robertson et al., 2017b; Sirks et al., 2024). Theoretical studies to model the offsets that arise from the self-interactions effectively decelerating the two DM halos when they pass through each other while the galaxies are only indirectly affected via gravity have been conducted to gain insights into their evolution and allow for constraints from observed systems to be inferred. Offsets between the galaxy and DM distributions are measured employing strong gravitational lensing to infer the total mass distribution. These measurements are challenging, and previous claims of large offsets have been questioned (e.g. Bradač et al., 2008; Dawson et al., 2012; Dawson, 2013; Jee et al., 2014, 2015; Harvey et al., 2017; Peel et al., 2017; Taylor et al., 2017; Wittman et al., 2018, 2023). Until today, observed systems show no clear deviation from collisionless DM, and mainly upper bounds on the self-interaction cross section from DM-galaxy offsets exist.
@@
This work is structured as follows. In Sect. 2, we first introduce the various exemplary differential cross sections considered in this work and review the rSIDM and fSIDM schemes for describing angle-dependent self-scattering. We introduce the new hSIDM scheme in Sect. 3 and furthermore discuss an analytically solvable test set-up that we use for validation later on. The numerical implementation is discussed in Sect. 4, and we present our validation checks in Sect. 5. The hSIDM scheme is used to simulate galaxy cluster mergers for SIDM with angle dependence as predicted by light mediator models in Sect. 6. Our main results for DM-galaxy offsets and their model-dependence is discussed in Sect. 7. Finally, we summarise and conclude in Sect. 8. Additional information can be found in the Appendices.
2 Review of angle dependence in self-interacting dark matter
@@
In this section, we first briefly review the angle dependence of typical differential cross sections relevant for light mediator models of SIDM as well as various definitions of angle-averaged cross sections considered in this context. We then review the existing rSIDM and fSIDM approaches for describing rare large-angle and frequent small-angle scatterings, respectively, on which the hybrid scheme introduced in this work is based.
@@ 2.1 Differential scattering cross sections
Light mediator models are characterised by scattering via a Yukawa interaction, for which DM particles of mass $m_{\chi}$ scatter via the exchange of a mediator with mass $m_{\phi}$. One can further discriminate the cases in which the two incoming DM particles are identical (we refer to this case as ‘Møller scattering’ in analogy to $e^{-}e^{-}$ scattering, noting however that the mediator particle is massive here), or distinguishable (being e.g. a particle and an antiparticle, referred to as ‘Rutherford scattering’, but again representing the case of a massive mediator in this work). In the non-relativistic limit and in Born approximation, Møller scattering receives contributions from $t$- and $u$-channel diagrams, yielding
@@
with the factor of one-half accounting for the double counting of the identical particles, while only the $t$-channel process contributes to Rutherford scattering, giving
@@
Following the notations used in Girmohanta & Shrock (2022); Robertson et al. (2017a), we introduced a parameter $\sigma_{0}$ that is related to the Yukawa coupling strength $\alpha_{\chi}$ via $\sigma_{0}=\alpha_{\chi}^{2}m_{\chi}^{2}/m_{\phi}^{4}$. More importantly, the angle dependence is related to the parameter
@@
where $\beta_{\mathrm{rel}}=v_{\mathrm{rel}}/c$ is the relative velocity of the incoming particles. The value of $r$ characterises the dependence of the cross section on the deflection angle $\theta$, with the isotropic limit being approached for $r\ll 1$, while scattering becomes strongly forward-dominated (Coulomb-like) for large values of $r$. For illustration, we show the angle dependence of the Møller cross section in Fig. 1 for $r=100$ and $10^{4}$, respectively, as well as an isotropic cross section for comparison. We emphasise the pronounced peaks towards small scattering angles (note the logarithmic $y$-axis). Since the Møller cross section is symmetric under $\theta\to\uppi-\theta$, it is also enhanced for $\theta\to\uppi$, due to the fact that both scattering particles are indistinguishable.
@@
Figure 1: Illustration of the differential cross section for self-scattering of identical particles via a light mediator (referred to as Møller scattering in this work) for two values of the anisotropy parameter $r$ defined in Eq. (3) (solid lines). Noting the logarithmic $y$-axis, the strong enhancement for $\theta\to 0$ and $\theta\to\uppi$ can be seen, and it becomes more pronounced the larger the $r$ (i.e. the lighter the mediator mass). We note that the isotropic case (red dashed) is not flat due to the scaling of the cross section with $\sin(\theta)$ (such that the lines represent the scattering rate within a given range $(\theta,\theta+\mathrm{d}\theta)$ of the azimuthal angle, for any value of the polar angle). We also illustrate the division into small and large-angle scattering by a critical angle $\theta_{\mathrm{c}}$, for two representative values, exemplifying the hybrid approach studied in this work.